OpenStudy (canimcan):

How can I finish this integral? My Result of integral: x(ln^2x+4lnx+1) is (3x^2lnx)+(x^2)/(2)-(x^3lnx)/(3)-(5x^3)/(9) But when I do the limit (from e^-2-sqrt3 to -2+sqrt3) I can't get result 123sqrt3 . What I am doing wrong?😕

OpenStudy (canimcan):

Can someone help?

OpenStudy (phi):

You could group your integral this way \[ \int (x \ln^2 x + x \ln x)+(2 x \ln x + x) + x \ln x \ dx\] the terms in parens are "exact" integrals (up to a factor): \[ \int \frac{1}{2}\left(2x \ln^2 x + 2x \ln x\right)+(2 x \ln x + x) + x \ln x \ dx\] i.e \[ \int \frac{1}{2} d(x^2 \ln^2 x) + d(x^2 \ln x) + x \ln x \ dx\] thus we get \[ \frac{1}{2} x^2 \ln^2 x + x^2 \ln x + \int x \ln x \ dx\] we can use integration by parts to evaluate the last integral (let u = ln x and dv = x dx) the last integral gives us \[ \frac{1}{2} x^2 \ln x - \frac{1}{4} x^2 \] combining like terms \[ \frac{1}{2} x^2 \ln^2 x + \frac{3}{2} x^2 \ln x- \frac{1}{4} x^2 \] However, I can't make out what the limits are

OpenStudy (canimcan):

Limits are from e^(-2-sqrt3) to (-2+sqrt3).

OpenStudy (holsteremission):

Not that this makes things much easier, but you could alternatively substitute \(\ln x=u\) so that \(e^u=x\) and \(e^u\,\mathrm{d}u=\mathrm{d}x\). Then \[\int_{e^{-2-\sqrt3}}^{e^{-2+\sqrt3}}x\left(\ln^2x+4\ln x+1\right)\,\mathrm{d}x=\int_{-2-\sqrt3}^{-2+\sqrt3} e^{2u}\left(u^2+4u+1\right)\,\mathrm{d}u\](... if only to clean up those limits. I'm also assuming you meant the upper limit to be the argument to \(\exp\)?) The resulting integral is, like its previous form, a perfect candidate for integration by parts.